### v2

# A Hierarchical Approach to Quantum Many-Body Systems in Structured Environments

Abstract
Open quantum systems that feature non-Markovian dynamics are routinely solved using techniques such as the Hierarchical Equations of Motion (HEOM). However, their usage of the entire system density-matrix renders them intractable for many-body systems. Here, we combine the HEOM with the Bogoliubov–Born–Green–Kirkwood–Yvon (BBGKY) hierarchy to achieve a rigorous description of open many-body systems in contact with structured photonic and phononic baths. We first rationalize that this stacked hierarchy accounts for spin-squeezing and superradiant emission despite its applicability to arbitrarily many emitters. The full potential of BBGKY-HEOM is then illustrated for two relevant applications: (i) the explicit treatment of vibrational modes provides access to resonant enhancements in few-emitter lasing, and (ii) the impact of phononic coupling and charge noise on many-body electronic systems embedded in host materials (e.g. molecules in organic crystals) is as relevant as electronic correlation. Our work establishes an accessible, yet rigorous, route between condensed matter and quantum optics, fostering the growth of a new domain at their interface.
I Introduction
The ability to control many-body quantum systems with quantized light is imperative in modern quantum technologies, including photonics, quantum memories, quantum sensing, or quantum information.[1, 2, 3, 4] An idealized description of the fundamental building block, the quantum emitter, is challenged by disturbances, such as charge noise or phonon scattering, that hint at the inherent many-body character of each quantum system. Strong light-matter coupling allows to control interactions over extended length scales[5, 6, 7, 8, 9], holding promise to design materials [10, 11, 12, 13, 14, 15], devices [16, 17, 18], or even chemical properties [19, 20, 21, 22, 23] non-intrinsically and on demand. The many-body physics originating from electronic and vibrational structure, as well as their respective decoherence, is crucial for a comprehensive understanding and refinement of this control-strategy.[24, 25, 26, 27, 28] Relevant experimental realizations include molecules or solids in micro- and plasmonic cavities, requiring a holistic treatment of electronic many-body physics coupled to various optical and phononic modes.[29, 16]
Our primary challenge in this study is to address physical processes wherein many-body interactions, strong light-matter coupling, and strong system-environment coupling are all fundamental. Achieving this goal necessitates combining concepts from quantum optics and solid-state physics. Quantum optics often treats light-matter interaction non-perturbatively and offers many sophisticated techniques for dealing with strongly coupled environments [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], but it typically simplifies the structure of the materials drastically. Conversely, condensed matter or quantum chemistry approaches accurately describe the material system but oversimplify interactions with surrounding environments.[47, 11, 12] Here, we show how to bridge system-bath, quantum optical, and condensed matter approaches. We establish a framework for studying the effects of strongly coupled bosonic fields on many-body physics more efficiently than previously possible.
In order to deal with the strongly coupled environments we rely on the Hierarchical Equations Of Motion (HEOM) [45, 46] approach. HEOM has been used successfully to solve a large variety of problems [48, 49, 50, 51, 52, 53, 54, 55, 56]. It lends itself to our cause particularly well, because (i) its approximation of the environment is completely independent of the system dimension, and (ii) it can be modified to incorporate nonlinear terms (see Sec. II.2). However, it treats the system part exactly and thus fails for many-body systems due to the exponential scaling of the system dimension.
Wavefunction-based approaches, such as tensor networks [57] or matrix product states, provide in principle exact and reliable results [58, 59, 60], but increase exponentially in cost over time. Their application is practically limited to short time scales of low-dimensional systems. Non-equilibrium Green functions approaches [61] made recent progress after the successful integration of the generalized Kadanoff-Baym ansatz [62, 63, 64] but a consistent treatment of complex structured baths remains challenging. An attractive approach with good scaling properties is to perform a cluster expansion of the density matrix into single-, two-, three-, and $N$-body reduced density matrices and then truncate the resulting Bogoliubov–Born–Green–Kirkwood–Yvon (BBGKY) hierarchy at some low order [65, 66].
Here, we combine BBGKY and HEOM to establish a stacked BBGKY-HEOM hierarchy that provides an efficient framework for the simulation of correlated many-body systems subject to structured baths. We apply our BBGKY-HEOM method to an experimentally relevant few-emitter lasing setup comprising multiple molecules inside a plasmonic cavity. There, we explicitly treat the vibrational modes of each molecule in a realistic manner by using parameters obtained from density-functional theory (DFT) calculations. Our results go beyond the state of the art and give new insights into how the vibrational spectrum influences the lasing process. Subsequently, we go past the paradigm of the two-level emitter and demonstrate that BBGKY-HEOM is able to capture many-body electronic dynamic in the Fermi-Hubbard model coupled to structured environments. BBGKY-HEOM provides accurate predictions as long as the onsite repulsion $U$, and with it the degree of electronic correlation, remains moderate. We notice that the bath adds a stabilizing contribution to the BBGKY equations.
The remainder of this article is structured as follows: Sec. II provides a short introduction to non-Markovian quantum systems and many-body theory before the derivation of the BBGKY-HEOM equations in Sec. II.2. We illustrate their performance, limitation, and associated physical intuition for various systems in Sec. III. After benchmarking our method for many-emitter quantum optics in Sec. III.1, we showcase its potential by capturing vibrationally enhanced few-emitter lasing in Sec. III.2. We subsequently shift our focus to many-body electronic systems in Sec. III.3, where we investigate the Fermi-Hubbard model coupled to a lossy cavity mode and subject to charge noise . This system in then embedded into a highly structured bath in Sec. III.4 that represents phononic motion in organic crystals. Sec. IV finally concludes our study and provides an outlook into the near future.
II Theory
A material comprising electrons and nuclei moving at non-relativistic velocities is described by Schrödinger’s equation. Its interplay with electromagnetic fields necessitates the consideration of the (quantized) normal modes of Maxwell’s equations. This combined system can be described using the minimal coupling Hamiltonian. In Coulomb gauge,
[…]
for $N_{e}$ electrons and $N_{n}$ nuclei with charge $q_{i}$ and mass $m_{i}$. We will use atomic units in the following. Eq. (1) splits the transversal degrees from the longitudinal and instantaneous Coulombic interactions $H_{\parallel}$. The latter includes electronic and nuclear interactions which are initially treated in first quantization
$\begin{split}\hat{H}_{\parallel}&=\hat{V}_{ee}+\hat{V}_{en}+\hat{V}_{nn};\quad\hat{V}_{ee}=\sum_{i\neq j}^{N_{e}}\frac{1}{|\textbf{r}_{i}-\textbf{r}_{j}|}\\ \hat{V}_{en}&=\sum_{i,j}^{N_{e},N_{n}}\frac{-q_{i}}{|\textbf{r}_{i}-\textbf{R}_{j}|};\quad\hat{V}_{nn}=\sum_{i\neq j}^{N_{n}}\frac{q_{i}q_{j}}{|\textbf{R}_{i}-\textbf{R}_{j}|}.\end{split}$ (2)
Their interplay results in rich many-body physics that gives rise to electronic, vibronic, and phononic structure and, unfortunately, to an exponentially increasing complexity. The open quantum systems approach now aims to identify a small subsystem of interest that we can treat explicitly (the ”system”) while absorbing the remainder into environments (the ”baths”) [67]. The first step is to simplify the Hamiltonian to a form that contains only linear interactions between system and bath(s)
$\hat{H}_{tot}=\hat{H}_{sys}+\hat{H}_{B}+\hat{H}_{sys,B}.$ (3)
The gauge freedom of the electromagnetic modes provides various different paths for this. An intuitive approach is to isolate the bilinear term $\sum_{i}-i\hbar q_{i}/m_{i}\hat{\textbf{A}}(\textbf{r}_{i})\cdot\bm{\nabla}_{i}$ as coupling operator and absorb the quadratic diamagnetic term into dressed photonic operators [68, 69]. A second direction, often more promising for finite systems, is to shift into the multi-polar gauge (detailed in App. E). The combination with the long-wavelength approximation then provides a simple bilinear coupling between dipole moment and displacement field. In addition, self-polarization terms modify the dynamics of the system and the redefinition of the optical creation and annihilation operators introduce the need to reconsider the interpretation of system and bath. Vibrational and phononic interaction follow similarly. The influence of nuclear motion on the electron-nuclear interaction can be expanded around the equilibrium configuration
$\hat{V}_{en}\approx\hat{V}_{en}|_{0}+\sum_{j}\delta\textbf{R}_{j}\cdot\bm{\nabla}_{j}\hat{V}_{en}|_{0}+...$ (4)
and truncated at harmonic order, resulting in a bilinearly coupled electron-phonon or electron-vibration coupling that depends on the electronic configuration. Note that also here second-order (known as Debye-Waller) corrections appear that require to carefully scrutinize the separation between system and bath [70]. The following derivations and demonstrations are agnostic to such subtleties, as we will focus on the development and illustration of the BBGKY-HEOM methodology. A thorough derivation of system-bath couplings for a specific physical realization is evidently essential for meaningful predictions but it is not the focus of this manuscript.
II.1 Many-body Open Quantum Systems beyond the Markovian Limit
Following the steps outlined above, both the phonon and the (transversal) photon sector of $\hat{H}_{mc}$ can be described as a collection of quantum harmonic oscillators that couple linearly to the electronic system. This description of an environment is frequently the case in quantum optics [71], or when the system is coupled weakly to a large number of degrees of freedom where central limit type of arguments yield an approximately Gaussian response [72]. We label these distinct environments with $k$, such that the bosonic annihilation (creation) operators belonging to mode $\lambda$ of environment $k$ are given by $a_{\lambda}^{k}$ ($a_{\lambda}^{k\dagger}$). The Hamiltonian corresponding to the total state $\rho_{tot}$ of system and environments is then
$H=$ $H_{sys}+\sum_{k,\lambda}g_{\lambda}^{k}(L^{k}a_{\lambda}^{k,\dagger}+L^{k\dagger}a^{k}_{\lambda})+\sum_{\lambda}\omega^{k}_{\lambda}a_{\lambda}^{k\dagger}a^{k}_{\lambda}.$ (5)
Here $L^{k}$ is the system operator that describes the coupling to the non-Markovian environment $k$ and $\omega_{\lambda}^{k}$, $g_{\lambda}^{k}$ are the frequencies and coupling strengths of the respective modes. The environments are characterized by their spectral density $J^{k}(\omega)=\sum_{\lambda}\lvert g^{k}_{\lambda}\rvert^{2}\delta(\omega-\omega^{k}_{\lambda})$ or equivalently by their bath correlation function $\alpha^{k}(\tau)=\int_{0}^{\infty}J_{k}(\omega)e^{-i\omega\tau}\mathrm{d}{\tau}$. When a description of the environment in terms of harmonic oscillators is suitable and the inital state of system and environment is a product state, the HEOM approach yields a numerically exact description for the evolution of the open quantum system, even for strong coupling [45, 46, 48]. To use HEOM, the bath correlation function needs to be fitted with exponential functions [73] according to
$\alpha^{k}(\tau)\approx\sum_{j}^{N_{exp}}G_{j}^{k}\exp\left(-w_{j}^{k}\tau\right),$ (6)
with $G_{j}^{k},\;w_{j}^{k}\in\mathbb{C}$. The resulting hierarchy couples the physical density matrix to auxiliary density matrices, and the number of auxiliary density matrices depends on the number of exponentials $N_{exp}$ as well as the system-environment coupling strength. In fact, HEOM type of techniques can be even extended to cases where the environment is anharmonic [74, 75]. In this work, we focus on the case of harmonic baths, either corresponding to optical modes confined in an imperfect cavity or a phonon bath into which a quantum emitter is embedded. The influence functional of the bath can then be computed exactly under the harmonic assumption by using the path integral formalism [76]. HEOM follows from this influence functional [46] and we provide a sketch of the derivation for the open system in App. B.
Following the previous motivation, we start with a many-body system that is linearly coupled to a collection of harmonic oscillators. The $N$-body system can be decomposed into a single particle Hamiltonian $H_{i}$ and two-particle interactions $V_{ij}$
$\begin{split}H_{sys}=&\sum_{i=1}^{N}H_{i}+\sum_{i\neq j}^{N}V_{ij}.\end{split}$ (7)
Additional (local) Markovian dissipative processes of the system are captured by $\mathcal{L}_{diss}$. The evolution of the total state $\rho_{tot}$ of system and environments is then described by
$\dot{\rho}_{tot}=$ $-i[H,\rho_{tot}]+\mathcal{L}_{diss}(\rho_{tot}).$ (8)
Eq. (8) together with Eq. (5)(7) provide the general form for systems discussed in this work. HEOM retains information about the bath, allowing one to access e.g. photonic correlation functions.
For the coupling operators $L_{k}$ we distinguish two cases, both illustrated in the following. First, all particles couple to the same environment with equal strength and the coupling operators take the form
$L^{k}=\sum_{i=1}^{N}L^{k}_{i}.$ (9)
Second, each particle couples to its own local environment, however, all local environments are identical. This situation arises, for example, in a collection of molecules, each with its local vibrational bath (see Sec. III.2). The coupling operators are then denoted as $S_{k}$, and act only on the local Hilbert space of particle $k$. The evolution equation (5) is invariant under particle exchange. Such assumptions are natural for indistinguishable particles but one can relax this condition if one would like to describe ensembles of inhomogeneously broadened emitters.
Despite the restrictions employed above, the resulting system-bath Hamiltonian applies to a vast range of systems critical in fields such as quantum optics, quantum impurities in solid hosts, and strongly correlated materials in cavities.
II.2 Combining HEOM with BBGKY
Let us now tackle the exponential growth of the Hilbert space by stacking HEOM and BBGKY hierarchy. We focus here on a single environment for the sake of simplicity, but note that the extension to multiple environments is trivial and will be illustrated in Sec. III.4. The HEOM equation accounts for the non-Markovianity of the coupled environment with an infinite set of coupled equations for the matrices $\rho^{(\bm{n},\bm{m})}$, whose dimension is equal to the dimension of the system Hilbert space. Here, $\rho^{(\bm{0},\bm{0})}=\operatorname{Tr}_{B}(\rho_{tot})=:\rho_{sys}$ is the physical density matrix of the system and $\bm{m},\bm{n}$ are vector indices that label auxiliary density matrices. The hierarchy is truncated at a sufficiently high order which is determined by the non-Markovianity of the system-bath interaction. A Markovian bath corresponds to a truncation at first order according to $\rho^{(\bm{e}_{k},\bm{0})}\propto L\rho^{(\bm{0},\bm{0})}$, $\rho^{(\bm{0},\bm{e}_{k})}\propto\rho^{(\bm{0},\bm{0})}L^{\dagger}$, where $\bm{e}_{k}$ denotes the unit vector in direction $k$. The evolution equation for the auxiliary density matrices (including the physical reduced density matrix $\rho^{(\bm{0},\bm{0})}$) for a single environment with coupling operator $L$ reads
$\dot{\rho}^{(\bm{n},\bm{m})}$ $=-i[H_{sys},\rho^{(\bm{n},\bm{m})}]-\left(\bm{w}\cdot\bm{n}+\bm{w}^{*}\cdot\bm{m}\right)\rho^{(\bm{n},\bm{m})}$
$+\sum_{j=1}^{N_{exp}}\bigg{(}G_{j}n_{j}L\rho^{(\bm{n}-\bm{e}_{j},\bm{m})}+G_{j}^{*}m_{j}\rho^{(\bm{n},\bm{m}-\bm{e}_{j})}L^{\dagger}$
$\quad\quad\quad+\left[\rho^{(\bm{n}+\bm{e}_{j},\bm{m})},L^{\dagger}\right]+\left[L,\rho^{(\bm{n},\bm{m}+\bm{e}_{j})}\right]\bigg{)}.$ (10)
Bold symbols indicate the vector character of a variable and the complex coefficients $G_{j},\bm{w}$ can be obtained from a fit according to Eq. (6) or inferred from scattering theory [77]. The application to many-body systems now requires us to identify a path to limit the dimensionality of the descriptor, so far $\rho^{(\bm{n},\bm{m})}$, to manageable levels.
An intuitive way to accomplish this is presented by the BBGKY hierarchy which is build around the reduced density matrices (RDM)
$\begin{split}F_{123}^{(\bm{n},\bm{m})}=&N(N-1)(N-2)\operatorname{Tr}_{4,5,...,N}(\rho^{(\bm{n},\bm{m})}),\\ F_{12}^{(\bm{n},\bm{m})}=&N(N-1)\operatorname{Tr}_{3,4,...,N}(\rho^{(\bm{n},\bm{m})}),\\ F_{1}^{(\bm{n},\bm{m})}=&N\operatorname{Tr}_{2,3,...,N}(\rho^{(\bm{n},\bm{m})}).\end{split}$ (11)
Here $\operatorname{Tr}_{i,i+1,...,N}$ denotes the trace over all particles with index larger than $i$. The reduced density matrices are normalized to account for the particle number under contraction [78]. Note that since we are restricting ourselves to permutationally symmetric systems, the order of indexing is irrelevant.
Our goal in the following is to derive a closed hierarchy of equations for the reduced density matrix $F_{12}$, such that the (auxiliary) density matrices in Eq. (II.2) only scale with the square of the dimension of the single-particle Hilbert space. From Eq. (II.2) it follows that (see App. C)
$\dot{F}_{12}^{(\bm{n},\bm{m})}$ $=-i[H_{1}+H_{2},F_{12}^{(\bm{n},\bm{m})}]-i\operatorname{Tr}_{3}\left([V_{13}+V_{23},F_{123}^{(\bm{n},\bm{m})}]\right)-\left(\bm{w}\cdot\bm{n}+\bm{w}^{*}\cdot\bm{m}\right)F_{12}^{(\bm{n},\bm{m})}$
$+\sum_{k=1}^{M}\bigg{(}G_{k}\left(n_{k}(L_{1}+L_{2})F_{12}^{(\bm{n}-\bm{e}_{k},\bm{m})}+n_{k}\operatorname{Tr}_{3}(L_{3}F_{123}^{(\bm{n}-\bm{e}_{k},\bm{m})})\right)+G_{k}^{*}\left(m_{k}F_{12}^{(\bm{n},\bm{m}-\bm{e}_{k})}(L_{1}^{\dagger}+L_{2}^{\dagger})+m_{k}\operatorname{Tr}_{3}(L_{3}^{\dagger}F_{123}^{(\bm{n},\bm{m}-\bm{e}_{k})})\right)$
$\quad\quad+\left[F_{12}^{(\bm{n}+\bm{e}_{k},\bm{m})},L_{1}^{\dagger}+L_{2}^{\dagger}\right]+\left[L_{1}+L_{2},F_{12}^{(\bm{n},\bm{m}+\bm{e}_{k})}\right]\bigg{)}.$ (12)
For any interacting many-body system, the two-body RDM $F_{12}$ will couple to the three-body RDM $F_{123}$, which in turn will couple to the four-body RDM and so on. In order to close this hierarchy we express the three-body RDM in terms of the one- and two-body RDM
$F_{123}\approx\widetilde{F}_{123}(F_{12},F_{1}),$
where $\widetilde{F}_{123}$ should be understood as a function that approximates the true three-body RDM. We demonstrate in App. C how this can be achieved by reformulating the hierarchy of equations for the density matrices $\rho_{sys}^{(\bm{n},\bm{m})}$ as an equation for a single operator in an extended Hilbert space $\mathcal{H}_{ext}=\mathcal{H}_{sys}\otimes\mathcal{H}_{aux}$ and subsequently neglecting three-body correlations within the system. However, we keep correlations between two particles in the system and the auxiliary degree of freedom that captures the influence of the environment. With these approximations we arrive at the expression below for the three-body RDM, where we have left out indices $(\bm{0},\bm{0})$ for brevity
$\widetilde{F}^{(\bm{n},\bm{m})}_{123}$ $=4\frac{(N-1)(N-2)}{N^{3}}\operatorname{Tr}(F_{1}^{(\bm{n},\bm{m})})F_{1}F_{2}F_{3}$
$+\frac{N-2}{N}\left(F_{12}F_{3}^{(\bm{n},\bm{m})}+F_{2}^{(\bm{n},\bm{m})}F_{13}+F_{1}^{(\bm{n},\bm{m})}F_{23}\right)$
$+\frac{N-2}{N}\left(F_{1}F_{23}^{(\bm{n},\bm{m})}+F_{13}^{(\bm{n},\bm{m})}F_{2}+F_{12}^{(\bm{n},\bm{m})}F_{3}\right)$
$-\frac{N-2}{N^{2}}\operatorname{Tr}(F_{1}^{(\bm{n},\bm{m})})\left(F_{12}F_{3}+F_{13}F_{2}+F_{1}F_{23}\right)$
$-2\frac{(N-2)(N-1)}{N^{2}}\Big{(}F_{1}^{(\bm{n},\bm{m})}F_{2}F_{3}$
$\quad+F_{1}F_{2}^{(\bm{n},\bm{m})}F_{3}+F_{1}+F_{2}F_{3}^{(\bm{n},\bm{m})}\Big{)}.$ (13)
Inserting the above approximation into Eq. (II.2) provides the BBGKY-HEOM equation up to three-body correlations – the central tool of this work. As desired, it represents a closed hierarchy for the two-particle RDM that can account for interactions within the system and simultaneously includes the influence of a generic non-Markovian environment.
$\widetilde{F}^{(\bm{m},\bm{n})}_{123}$ is invariant under particle exchange, but does not preserve the (anti-)symmetry of the wavefunctions as required for (fermionic) bosonic particles. As explained in detail in App. C this can be achieved by employing an (anti-) symmetrization operator which projects $\widetilde{F}^{(\bm{m},\bm{n})}_{123}$ onto the (anti-) symmetric subspace of the Hilbert space while retaining its norm [66].
III Results
Every quantum system is a collection of individual particles, may it be a collection of single-photon emitters or the individual electronic and vibrational degrees of freedom that constitute one such emitter. In the following, we demonstrate the applicability, strength, and limitation of BBGKY-HEOM for both of these cases. In particular, we focus on three systems from quantum optics and condensed matter physics. Benchmarks via driven Tavis-Cummings and Cavity-Fermi-Hubbard models serve to demonstrate the excellent performance of BBGKY-HEOM. We then go beyond the state of the art with applications to few-emitter molecular lasing and system-bath competition in the Cavity-Fermi-Hubbard model.
Figure 1: Squeezing and superradiance in the driven Tavis-Cummings model: (a) Spin dynamics for $N=50$, $g\sqrt{N}/\kappa=0.5$, $\Delta_{z}/\kappa=0.2$, $\Omega/\kappa=0.05$, $\Delta_{c}/\kappa=10$. (b) Corresponding spin-squeezing (see text). (c) Superradiant emission from fully excited state. Parameters adjusted to $\Delta_{z}/\kappa=2$, $\Delta_{x}/\kappa=1$, $\Delta_{c}/\kappa=2.4$, $g/\kappa=0.1$. (d) Photon-occupation normalized by $N^{2}$ over time and maximum photon number in burst (not normalized) for different N obtained with BBGKY-HEOM. BBGKY-HEOM provides reliable predictions and only begins to deviate slightly at long times when explicit 3-body correlations accumulate to noticeable values.
III.1 Driven-dissipative many-emitter systems
A collection of emitters that share a common mode can interact with it collectively, potentially resulting in burst-like correlated emission known as superradiance and non-classical squeezed spin states. Following a series of simplifications for matter, cavity, and light-matter coupling, one obtains the widely used Dicke Hamiltonian [71]
$H_{Dicke}=\omega_{a}\sum_{i}\sigma_{z}^{i}+\omega_{c}a^{\dagger}a+g\sum_{i}\sigma_{x}^{i}(a^{\dagger}+a).$ (14)
The Pauli matrices $\sigma_{x,y,z}^{i}$ represent the SU(2) algebra of an idealized two-level emitter while $a^{\dagger}$($a$) are the creation (annihilation) operators of the cavity mode. The collection of emitters can be brought out of equilibrium by coherently driving the system $H_{drive}=2\Omega\cos{\omega_{d}t}\sum_{i}\sigma_{x}^{i}$. A unitary transformation moves us into a frame that rotates with the driving frequency $\omega_{d}$. As long as the drive is close to resonance, i.e., $\omega_{d}\gg\Delta_{z},\Delta_{x},g,\Omega$ with the detunings $\Delta_{z}=\omega_{a}-\omega_{d}$, $\Delta_{c}=\omega_{c}-\omega_{d}$, we can safely discard counter-rotating terms $e^{i(\omega_{d}+\omega_{a/c})}$ that oscillate too fast to influence the dynamics of the system. We obtain the paradigmatic driven Tavis-Cummings Hamiltonian
$\begin{split}H_{DTC}=&\Delta_{z}\sum_{i}\sigma_{z}^{i}+\Omega\sum_{i}\sigma_{x}^{i}+\Delta_{c}a^{\dagger}a\\ +&{g}\sum_{i}\left(\sigma_{+}^{i}a+\sigma_{-}^{i}a^{\dagger}\right).\end{split}$ (15)
We will consider the emitters or spins in the following as our system of interest and the cavity as an environment that dissipates energy into free space at a rate $\kappa$. The bath correlation function is readily obtained, for example, from the Heisenberg-Langevin equations [67], as
$\alpha(\tau)=g^{2}\langle a^{\dagger}(\tau)a(0)\rangle=g^{2}e^{(-i\Delta_{c}-\kappa)\tau}.$ (16)
[…]
Given the necessary computational resources, we could have just as well decided to directly solve the entire system of $N$-emitters plus lossy cavity using the Lindblad master equation
$\dot{\rho}=-i[H_{DTC},\rho]+\kappa\left(a\rho a^{\dagger}-a^{\dagger}a\rho-\rho a^{\dagger}a\right).$ (17)
This exact solution will be used to estimate the quality of the BBGKY-HEOM and can be obtained in a collective spin picture $S_{k}=\sum_{i}\sigma_{k}^{i}$, $k\in\{x,y,z,+,-\}$. The $2^{N}$ dimensional system Hilbert space can then be restricted to a basis of $N+1$ permutationally symmetric Dicke states {|$\{\ket{N/2,m},m=-N/2,-N/2+1,\dots,N/2\}$ which results in a linear increase of the relevant basis size with $N$ and therefore a quadratic increase of the Liouville space. The collective spin-picture fails as soon as individual emitter-emitter interactions need to be included, resulting in an exponentially growing Hilbert space that prohibits exact solutions for systems larger than a handful of emitters. However, the computational cost for BBGKY-HEOM is unaffected by this change and remains system-size independent. Even if there is no direct interaction in this model, the spins still interact indirectly via the cavity mode.
Fig. 1 (a) illustrates the system dynamics of Eq. (17) for $N=50$ emitters strongly coupled to the cavity $\sqrt{N}g/\kappa=\sqrt{N/4}$ in comparison to the exact solution of the master equation. The spin components $\langle S_{i}\rangle$ are accurately predicted within the simulation window, despite the non-classical dynamics of the system. Shown in the upper inset is the spin squeezing (see App. A) which is exhibiting values smaller than unity – a sufficient condition for entanglement. We observe small deviations at long times which can be understood by investigating the slow increase in the explicit 3-body correlations (lower inset), i.e., the term that is discarded in the currently chosen truncation. The spin-squeezing is further illustrated by the spin-Q function $Q=\bra{\theta,\phi}\rho\ket{\theta,\phi}$ in the upper right subplot (b), with |$\ket{\theta,\phi}$ the spin coherent state. Mean-field theory will fail in predicting any of the spin-squeezing effects.
Let us fully excite all emitters and illustrate the resulting superradiant emission in Fig. 1 (c) (parameters adjusted, see caption). BBGKY-HEOM is accurate, with small deviations to the exact solution only at long times near the fully de-excited steady-state. While mean-field theory models the initial frequency quite accurately, it fails to model the spontaneous emission process and remains trapped at the upper fix-point. The inset projects this dynamic on Bloch’s sphere. The failure of mean field originates from restricting the dynamics to the surface of the sphere, whereas the correct dynamics traverses the Bloch ball diagonally. The sudden increase in the photon number originating from the burst-like emission of $N$ correlated emitter is presented in Fig. 1 (d), demonstrating again accurate predictions by BBGKY-HEOM. Continuous driving results in a continuous replenishment of the cavity mode. The maximum photon-number (inset) is correctly described in its quadratic increase with the emitter number $N$. In conclusion, BBGKY-HEOM provides accurate predictions for matter and cavity observables at constant cost, i.e., for arbitrary many emitter and with the option to account for direct interaction.
III.2 Few-emitter lasing
Figure 2: Overview of the few emitter lasing system.(a) Sketch of the plasmonic nanocavity, comprised of a gold nanoparticle above a gold surface. (b) We restrict our model of the electronic states to the first bright transition, leading to a two level system. However, we take the full spectrum of vibrational modes into account, which we approximate as a large set of harmonic oscillators (purple). Each oscillator in turn couples to a continuum of vibrational modes that we model with an ohmic spectral density (blue). All the vibrations can be combined into a single effective bath with the spectral density $J_{eff}(\omega)$ shown in (d) (green solid line). The corresponding bath correlation function shown in the inset is then fitted with exponentials (black dahed line). Graphic (c) shows a sketch of the lasing mechanism.
For our next application, we turn to a system where no benchmark calculations are available and where BBGKY-HEOM delivers new insights for systems of current experimental relevance. As shown in Fig. 2 (a) a few ($N\approx 2-20$) methylene-blue (MB) molecules are placed inside a plasmonic nanocavity, consisting of an Au nanoparticle on a mirror [79, 80]. Each molecule has a complex internal structure which, in addition to the electronic degrees of freedom, consists of hundreds of internal vibrational modes that in turn couple to a continuum of vibrations in the surrounding dielectric medium. Reference [79] treated this system with the help of simplified effective models, which provided valuable analytic insights and qualitatively captured many experimental features. To achieve these analytical insights, the structure of the vibrational spectrum was simplified to an incoherent drive – a widely applied approximation [81]. However, such simplifications ignore an integral part of molecular physics, limiting understanding and agreement with experiment. We demonstrate in the following how our BBGKY-HEOM method can be used to take the next step and obtain much more detailed results by retaining the vibrational structure. Our approach reveals additional resonant features that are not visible in paradigmatic models.
We restrict the electronic degrees of freedom for each molecule to the many-body states associated with the first bright electronic transition $\widetilde{\omega}_{0}$. The full spectrum of vibrational modes with frequencies $\omega^{vib}_{\lambda}$ and Huang-Rhys factors $S_{\lambda}$ are taken into account. Frequencies and Huang-Rhys factors are estimated using time-dependent density-functional theory (see App. F.1). In the double-harmonic approximation for the electron-phonon coupling, each of the 107 vibrational normal modes is described as a harmonic oscillator, where the equilibrium position of the oscillator is shifted by a displacement $\Delta_{\lambda}=\sqrt{2S_{\lambda}/\omega_{\lambda}}$ if the molecule is in the electronically excited state. Within this approximation, our Hamiltonian for a single molecule takes the form
$H_{mol}=$ $\frac{\widetilde{\omega}_{0}}{2}\sigma^{z}+\sum_{\lambda=1}^{107}\frac{p_{\lambda}^{2}}{2}+\frac{(\omega_{\lambda}^{vib})^{2}}{2}\left(q_{\lambda}-\Delta_{\lambda}\sigma_{+}\sigma_{-}\right)^{2},$
$=$ $\frac{\omega_{0}}{2}\sigma^{z}-\frac{1}{\sqrt{2}}\sum_{\lambda=1}^{107}{\omega_{\lambda}^{vib}}\sqrt{S_{\lambda}}\sigma_{+}\sigma_{-}(b_{\lambda}+b_{\lambda}^{\dagger})$ (18)
$+$ $\sum_{\lambda=1}^{107}\omega_{\lambda}^{vib}b_{\lambda}^{\dagger}b_{\lambda}$
Figure 3: Results for the few emitter lasing system. We compare the qualitative steady state behavior for increasing the coherent drive in the full model (III.2),(20) against an effective incoherent drive that neglects the vibrational structure (III.2). Plots (a) and (b) show the occupation of the excited electronic state, $p_{e}$, and the cavity mode, $a^{\dagger}a$, respectively, for a plasmonic cavity containing ten molecules. Both plots show maxima arising from resonances with the vibrational spectrum that are missed by an effective treatment with an incoherent drive. Those maxima become more pronounced for increasing $N$ as shown in plot (c).
where $q_{\lambda}$, $p_{\lambda}$ are the mass renormalized conjugate position and momenta of the $\lambda$-th normal mode (see Appendix F.1 for a more detailed derivation). The second line reformulated the Hamiltonian in terms of the annihilation (creation) operators $b_{\lambda}$ ($b_{\lambda}^{\dagger}$), with $b_{\lambda}=\sqrt{\omega_{\lambda}/2}(q_{\lambda}+ip_{\lambda}/\omega_{\lambda})$ and we introduced the effective excitation energy $\omega_{0}=\widetilde{\omega}_{0}+\sum_{\lambda}\omega_{\lambda}^{2}\Delta_{\lambda}^{2}/2$, here $\omega_{0}=20857\mathrm{cm}^{-1}$. Due to the embedding of the MB molecules in the surrounding dielectric host medium [80], each of these vibrational modes in turn couples to a continuum of environmental modes, which we describe with an ohmic spectral density (see Fig. 2 (b) and Appendix F.2). These environmental modes lead to a broadening of the vibrational spectrum and ensure that the vibrational modes quickly relax to the corresponding equilibrium position. Due to the rapid relaxation, the setup resembles a four-level lasing scheme in the sense of Fig. 2 (c). Coherent excitation to state |$\ket{4}$ is followed by a fast relaxation into the new equilibrium position of the vibrational modes |$\ket{3}$111 Such a two-step relaxation process of coherent driving combined with vibrational relaxation is often approximated as an effective incoherent drive [81, 79]. This eliminates the presence of vibrational modes, implying a form of wide-band approximation that disregards any structural features. , from which the state decays into |$\ket{2}$ via cavity emission. The decay is then again followed by a fast vibrational relaxation into |$\ket{1}$.
To apply BBGKY-HEOM we proceed by combining all 107 vibrational modes in an exact way into the single (local) bath with effective spectral density [83] $J_{eff}(\omega)=\sum_{\lambda}\lvert g^{eff}_{\lambda}\rvert^{2}\delta(\omega-\omega_{\lambda})$ shown in Fig. 2 (d). This process is described in detail in Appendix F.2.
A few ($N$) of those molecules are now driven by a coherent field with amplitude $2E_{d}$ and frequency $\omega_{0}$. In addition, each molecule is coupled equally with strength $g^{cav}$ to a cavity in resonance with $\omega_{0}$. As detailed in Appendix F.3 we can go into an interaction picture that rotates with the drive frequency $\omega_{0}$ and perform a rotating wave approximation, assuming that $\omega_{0}\gg E_{d},g^{cav}$. The resulting Hamiltonian in the interaction picture then takes the form of Eq. (5):
$H_{tot}^{I}$ $=E_{d}\sum_{j}\sigma_{x}^{j}$
$-\sum_{j}\left(\sum_{\lambda}g_{\lambda}^{eff}\sigma_{+}^{j}\sigma_{-}^{j}(b_{\lambda,j}^{\dagger}+b_{\lambda,j})+\sum_{\lambda}\omega_{\lambda}^{eff}b_{\lambda,j}^{\dagger}b_{\lambda,j}\right)$
$+\sum_{\lambda}g_{\lambda}^{cav}\left(\sum_{j}\sigma_{+}^{j}a_{\lambda}+\sum_{j}\sigma_{-}^{j}a_{\lambda}^{\dagger}\right)$
$+\sum_{\lambda}(\omega_{\lambda}^{cav}-\omega_{0})a_{\lambda}^{\dagger}a_{\lambda}.$ (19)
The second line describes the coupling of each molecule to its local vibrational bath with spectral density $J_{eff}(\omega)$ and the third line describes the coupling to the (global) cavity environment with the Lorentzian spectral density $J_{cav}$. Additionally, we include spontaneous emission according to
$\dot{\rho}_{tot}=-i[H_{tot}^{I},\rho_{tot}]+\Gamma_{\downarrow}\sum_{j}\mathcal{D}[\sigma_{-}^{j}](\rho_{tot}).$ (20)
To apply BBGKY-HEOM we now fit the bath correlation function $\alpha_{eff}(\tau)$ with a sum of 5 exponentials as $\alpha_{eff}(\tau)\approx\sum_{i}G_{i}\exp(-W_{i}\tau)$ as shown in Fig. 2 (d). As the correlation function of the cavity $\alpha_{cav}(\tau)=G_{c}e^{-\kappa\lvert\tau\rvert}$ is already exponential, no additional fit is required. We now propagate the state with different drive strengths $E_{d}<E_{max}=0.1\omega_{0}$ and cavity parameters $G_{c}=0.2E_{d}$, $\kappa=5E_{d}$.
The results are shown in Fig. 3 (a), (b), where we compare the occupation of the excited electronic state $p_{e}$ and the cavity in the steady state for ten molecules and different coherent driving strengths $E_{d}$ (orange circles). The results are contrasted with a second set of BBGKY-HEOM simulations (black crosses), where the coherent drive and the vibrational bath have been combined into an effective incoherent pumping, such that the molecules evolve according to
$\dot{\rho}_{incoh}=$ $-i[H_{incoh}^{I},\rho_{incoh}]+\Gamma_{\downarrow}\sum_{j}\mathcal{D}[\sigma_{-}^{j}](\rho_{incoh})$
$+E_{d}\sum_{j}\mathcal{D}[\sigma_{+}^{j}](\rho_{incoh}),$
$H_{incoh}^{I}=$ $\sum_{\lambda}g_{\lambda}^{cav}\left(\sum_{j}\sigma_{+}^{j}a_{\lambda}+\sum_{j}\sigma_{-}^{j}a_{\lambda}^{\dagger}\right)$
$+\sum_{\lambda}(\omega_{\lambda}^{cav}-\omega_{0})a_{\lambda}^{\dagger}a_{\lambda}.$ (21)
We focus on the qualitative behavior as coherent driving and incoherent pumping will result in different steady-states for a given value. Fig. 3 (a) demonstrates the required population inversion for a lasing transition. In contrast to the predictions from the incoherent drive, the full model does not indefinitely increase its inversion, but rather reaches a plateau at large driving strengths. There the coherent drive dominates the dynamics, leading to 0.5 inversion. Interestingly, we find a maximal inversion at $E_{d}\approx 0.3E_{max}$, with the corresponding time evolution of $p_{e}$ shown in the inset. For the steady state values of the cavity occupation in Fig. 3 (b) we find that the results for the full model feature distinct maxima on top of a background that looks similar to the results obtained from the incoherent drive. These maxima become even more pronounced for larger numbers of molecules $N$, as they are enhanced by the self-amplifying lasing process, see Fig. 3 (c).
Let us rationalize this effect by starting from a single dissipation-less vibrational mode for the molecule
$H_{simple}=E_{d}\sigma_{x}+g^{vib}\sigma_{+}\sigma_{-}(b+b^{\dagger})+\omega^{vib}b^{\dagger}b.$ (22)
Looking at Fig. 2 (c) we can see that the four level picture breaks down if the two equilibrium positions are so close that $\lvert\innerproduct{4}{3}\rvert\approx 1$. In that case the molecule would look more like a two-level system in which a population inversion is not possible by means of coherent driving. As $b+b^{\dagger}$ corresponds to the displacement from the ground state equilibrium position as defined in Eq. (III.2), a significant occupation of the vibrational modes is beneficial for the lasing process. In Appendix G we show that a resonance occurs if the driving strength $E_{d}$ is equal to $\omega^{vib}/2$, resulting in a maximized occupation of the vibrational mode. In fact, we find roughly $\langle b\rangle\propto 1/(\omega^{vib}-2E_{d})$. Thus, the displacement of the vibrational mode is enhanced at certain frequencies, leading to a more efficient lasing process as explained above. Indeed, comparing Fig. 3 (c) to Fig. 2 (d) we find that the maxima of $\langle a^{\dagger}a\rangle$ line up with the peaks in the spectral density according to $\omega^{peak}=2E_{d}^{peak}$. Furthermore, artificially removing the peak at $\omega\approx 1.3E_{max}$ in the spectral density (Fig. 2 (d)) we find that also the 2nd maxima in the cavity occupation vanishes, as shown in Appendix G.
The many-body picture employed in this section provides straight-forward access to vibronic coupling and recommends itself as a natural basis to describe few-emitter lasing. It should be noted that plasmonic surfaces are prone to strong nonradiative losses via Landau damping and subsequent hot-carrier generation, beneficially utilized in plasmonic catalysis [84, 85, 86]. Future work should therefore consider exploring the impact of electronic and vibrational heating on the optical characteristics of few-emitter lasing [87].
III.3 Many-body electronic systems coupled to non-Markovian optical environments
Figure 4: Fermi-Hubbard model coupled to lossy cavity mode: (Right) Dipole moment driven by time-dependent potential $V_{i}(t)=\frac{f(t)}{5(i+1)}$, where f(t) is a step function as shown by the grey dashed line and $U=0.1$. The example simulates charge noise due to the movement of electrons in the vicinity of the molecule. Inset illustrates onsite occupation.
We will now relax the simplification of the emitter and instead combine the dynamics of a correlated electronic system with a structured system-bath coupling. The electronic structure of most organic molecules is dominated by covalent bonds or conjugated $\pi$-systems, requiring many-body methods to faithfully model their electronic properties. An excitation in such a system cannot be represented by the excitation events of a single electron. Instead, the surrounding electronic cloud reacts and screens the polarization induced by the external field – every optical excitation in a molecule or solid is inherently a many-body process. Electronic correlation can reach substantial levels for e.g. transition metals, which contribute strongly localized $d$ and $f$ orbitals shielded from molecular orbitals. Moving two electrons to such a state costs a considerable amount of energy, as they strongly repel each other, resulting in electronic motion that is strongly correlated. A prototypical approach to describe such a system is given by the Hubbard model
$\begin{split}\hat{H}_{H}&=-J\sum_{\sigma}\sum_{i=1}^{N}\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i+1,\sigma}+\hat{c}_{i+1,\sigma}^{\dagger}\hat{c}_{i,\sigma}\\ &+U\sum_{i=1}^{N}\hat{c}_{i,+\frac{1}{2}}^{\dagger}\hat{c}_{i,-\frac{1}{2}}^{\dagger}\hat{c}_{i,-\frac{1}{2}}\hat{c}_{i,+\frac{1}{2}},\end{split}$ (23)
with the hopping rate $J$ between two localized Wannier orbitals and the onsite energy $U$. Usually $U>0$ as it costs energy to force two electrons on the same site. Weak onsite interaction $U/J\ll 1$ will only marginally limit the free hopping of electrons between the states while extreme values of $U/J\gg 1$ will result in states similar to Wigner crystals [88]. We focus here on small to moderate onsite interactions $U/J<1$ to account for the weak to moderate correlation in molecular systems.
Naturally, the electronic structure is embedded in an environment of optical (or vibro-phononic) modes, resulting in the emergence of polaritonic states if coupled sufficiently strong. Our BBGKY-HEOM approach allows to consistently describe and monitor how the dynamics of the lossy cavity mode imprints correlations in the electronic many-body system and vice versa. The corresponding minimal Hamiltonian for such a Cavity-Fermi-Hubbard system [89, 90]
$\begin{split}\hat{H}&=\hat{H}_{H}+\hbar\omega(\hat{a}^{\dagger}\hat{a}+\frac{1}{2})\\ &-\sqrt{\frac{\hbar\omega}{2\varepsilon_{0}V}}(\hat{a}+\hat{a}^{\dagger})\textbf{e}_{cav}\cdot\hat{\bm{\mu}}+\frac{1}{2\varepsilon_{0}V}(\textbf{e}_{cav}\cdot\hat{\bm{\mu}})^{2},\end{split}$ (24)
includes self-polarization contributions of the localized dipole moment $\hat{\bm{\mu}}=\textbf{e}_{chain}\sum_{\sigma}\sum_{i=1}^{N}r_{i}\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$. The local positions $r_{i}\in\{-\frac{N-1}{2}\Delta,...,0,...,+\frac{N-1}{2}\Delta\}$ as well as their spacing $\Delta$ are defined with respect to the center of the Power-Zienau-Wooley (PZW) gauge [68, 69]. We will use $J=1$, $\omega=0.5$, assign a cavity loss of $\kappa=0.2$, a coupling strength of $g=\sqrt{\frac{\hbar\omega}{2\varepsilon_{0}V}}=0.1$, and disregard the self-polarization for the moment. We emphasize that extensions to ultra-strong coupling [91] require the careful consideration of self-polarization terms [68, 69].
One strength of explicitly treating the many-body electronic system is that the electronic response to external stimuli, such as charge noise, can be explicitly modelled. Assume a short polyacetylene molecule with 4 sites hosting 4 electrons, represented by Eq. (24), and initially in equilibrium. A distortion in the surrounding structure suddenly traps an electron near our molecule, resulting in a repulsive Coulomb potential $V(t)=\sum_{i=0}^{3}V_{i}(t)c_{i}^{\dagger}{c}_{i}$ and $V_{i}(0^{+})=\frac{1}{5(i+1)}$. The subsequently induced dipole moment is presented in Fig. 4 (left) for three different values of onsite repulsion $U$. Without direct electronic interaction, the BBGKY-HEOM is nearly exact as the many-body correlation induced by the cavity is comparably small in this case. The quality of our approximation remains excellent for $U=0.1$ and slowly worsens for systems with sizeable correlation of $U=0.2$, as one would expect.
Charge will not only localize once but randomly tunnel in and out of the impurity. The dynamics originating from the random localization of charge on the dipole moment is shown in Fig. 4 (right) with the corresponding changes in occupation (inset) and time-dependent (gray dotted) potential. Clearly, hitting a resonance condition such as at $Jt=40$ can result in large deviations from the equilibrium state without noticeable influence on the quality of the BBGKY-HEOM predictions. One strength of the BBGKY-HEOM is therefore to resolve the many-body electronic dynamics that originates from time-dependent modulations which allows us to simulate e.g. charge noise, dynamics screening, or triplet-triplet annihilation explicitly – a feature that we plan to leverage further in the future.
Let us end this section by illustrating the limitations of the BBGKY hierarchy. The 4 sites of our Hubbard-chain host again 4 electrons but we force them initially to occupy the 1st and 3rd site only, i.e., our initial state has two doubly occupied and two empty sites. Naturally, this state is far from the correlated ground-state and results in a violent dynamics of density sloshing forward and backward. Fig. 5 illustrates the onsite occupation, i.e., the electronic density, of the initially occupied sites over time.
Figure 5: Strong dynamics in Cavity-Fermi-Hubbard model: Occupation for second and last site, both initially doubly occupied, over time. Sites 0 and 2 are initially empty, resulting in rapid charge-oscillations between the sites and a quick build-up of 3-body correlations. The BBGKY-HEOM becomes increasingly worse with ongoing time which is accompanied with the appearance of negative eigenvalues in the 1RDM that could potentially lead to instabilities.
The exact reference calculations (black dashed) follow our predictions closely up to $Jt\approx 20$ before the quality of the BBGKY-HEOM prediction deteriorates. We can relate this sudden loss of accuracy to the appearance of artificial negative eigenvalues in the reduced density matrix (see inset) – a failure caused by the fact that the BBGKY hierarchy is not contraction consistent by default. Contraction consistency refers to the fact that the 2RDM must be obtained from the trace of the approximated 3RDM. Neglecting this condition results in a build-up of negative eigenvalues that will inevitably lead to numerical instabilities. Fortunately, such effects can be circumvented by enforcing contraction consistency via more complex reconstruction functionals $\widetilde{F}_{123}$ as shown in [92, 66]. BBGKY-HEOM is remarkably stable when compared to the bare dynamics of the electronic system alone. We attribute this observation to the fact that the decoherence imprinted by the structured bath penalizes instabilities as large uncontrolled oscillations couple efficiently to the optical mode which dissipates a notable fraction of the energy and thus stabilized the dynamics of the system.
III.4 Many-body dynamics embedded into organic crystals
Figure 6: Cavity-Fermi-Hubbard model under the influence of a complex phonon bath: The evolution of the electronic dipole with (blue, solid) and without (gray, dashed) the phonon bath after a quench $V_{i}(0^{+})=\frac{4}{5(i+1)}$ induced by charge noise. Parameters are identical to Fermi Hubbard example in previous section with U=0.1. The inset shows the phonon spectral density with $\int J_{phon}(\omega)\mathrm{d}\omega=0.1^{2}$ (coupling identical to cavity mode) where the superohmic background is highlighted as black dotted line.
Many-body electronic systems, such as molecules or (an)organic frameworks, are often embedded in a liquid or crystalline environment. For example, organic molecules embedded in organic crystals can serve as a highly coherent emitter [1], quantum memory [93], or organic light-emitting diode. Their environment is a combination of local vibrational, extended phononic, and photonic modes, giving rise to a highly structured and complex spectral function. BBGKY-HEOM provides now the means to retain the complexity of the electronic structure and combine it with the full system-bath dynamic, i.e., extending beyond previous approaches that relied on a simplified description of the electronic system [94].
Fig. 6 demonstrates the impact of the non-Markovian phonon-bath on the dynamics of the dipole moment of the same 4-site Cavity-Fermi-Hubbard model ($U=0.1$) subsequent to a quench induced by a charge fluctuation $V_{i}(0^{+})=\frac{4}{5(i+1)}$. As the nuclei move closer together (further apart) their effective hopping rate increases (decreases), which we model by coupling the phonon-bath to the hopping operator:
$L^{phon}=\sum_{\sigma}\sum_{i=1}^{N}\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i+1,\sigma}+\hat{c}_{i+1,\sigma}^{\dagger}\hat{c}_{i,\sigma}.$ (25)
The corresponding phonon spectral density $J_{phon}(\omega)$ is illustrated in the inset, comprising a superohmic background that provides a simplified model for the linear dispersion of the acoustic phonon branch of the crystal as well as additional discrete modes. Neglecting the phonon bath (gray dashed) leads to a strong misrepresentation, which is much larger than the possible deviations caused by the BBGKY approximation (recall Fig. 4), and clearly shows why a holistic approach to the many-body dynamics of the system and the system-bath interaction is crucial for a predictive description.
IV Conclusion
No quantum system is perfectly isolated, and the description of its environment can be essential for our physical understanding. Confined optical or plasmonic modes and vibrational or phononic modes might even challenge the dynamics of a subsystem of interest by strongly coupling to the internal degrees of freedom. Energy deposited in the environment will then not simply decohere but act back on the system, resulting in non-Markovian, or memory dependent, system-bath interaction. The dynamics of the system has often been simplified to the degree that the many-body character inherited by every molecule or solid-state impurity has been lost.
In this study, we presented a rigorous hierarchical approach based on stacking of the reduced density-matrix (BBGKY) and system-bath (HEOM) hierarchies together. BBGKY-HEOM is flexible enough to tackle critical quantum optical and condensed matter problems, striking a balance between high accuracy and good performance. In particular, we demonstrated that correlated (superradiant) emission and spin squeezing, i.e., dynamics that extend beyond the capabilities of mean-field theory, are accurately captured.
Bolstered by this excellent performance, we move to describe few-emitter lasing of methylene-blue molecules in plasmonic nanocavities. Explicit treatment of the structured vibrational bath uncovers resonant enhancements in the lasing process that go beyond the paradigmatic incoherent drive model. To the best of our knowledge, this effect has not been discussed before. BBGKY-HEOM thus demonstrates its potential for a more detailed understanding of light-matter interaction in general, and few-emitter lasing in particular.
Many-body electronic systems, here in the form of the Fermi-Hubbard model, are critical in modern condensed matter theory. We illustrate the interplay between correlated system dynamics and structured baths by simulating different forms of charge noise. Again, excellent performance is observed as long as the onsite repulsion $U$ and the electronic dynamic remain moderate. Finally, we demonstrated the major strength of BBGKY-HEOM by simulating the interplay between the vibro-phononic bath of an organic crystal coupled to the many-body dynamics of the Cavity-Fermi-Hubbard model. Both contributions, i.e., non-Markovian system-bath coupling and many-body dynamics are equally important – clearly stressing the need for a more holistic perspective on solid-state quantum emitters.
Our hierarchical approach can be rigorously informed from first principles – bridging previously disconnected research fields. In this way, BBGKY-HEOM provides a powerful tool that can be used alongside established methods (like cumulant expansions, density-functional theory, etc.) to gather a more detailed and holistic understanding. Especially promising examples include molecules interacting with optical fields [95], may it be for quantum technological applications [1, 96], polariton condensation [97], or few-emitter lasing [79].
Acknowledgements.
We thank Walter Strunz, Alexander Eisfeld, and Göran Johansson for insightful discussions. C.S. acknowledges support from the Swedish Research Council through Grant No. 2016-06059 and funding from the Horizon Europe research and innovation program of the European Union under the Marie Skłodowska-Curie grant agreement no. 101065117. Partially funded by the European Union. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or REA. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Spin squeezing
The spin squeezing parameter is defined as $\xi^{2}(t)=N(\Delta S_{\textbf{n}_{1}})^{2}/(\langle S_{\textbf{n}_{2}}\rangle^{2}+\langle S_{\textbf{n}_{3}}\rangle^{2})$, where $\textbf{n}_{1}(t)\bot\langle\textbf{S}\rangle(t)$ is obtained, by calculating the vector perpendicular to the current spin direction that has the minimal associated variance $(\Delta S_{\textbf{n}_{1}})^{2}$.
Appendix B Derivation of the HEOM
We sketch here a derivation of the HEOM equations. We assume that the system is coupled to the environment via operator $q$, such that in an interaction picture with respect to the free Hamiltonian of the environment we have
$\begin{split}H_{tot}=&H_{sys}+q\left(B(t)+B^{\dagger}(t)\right),\\ B(t)=&\sum_{\lambda}g_{\lambda}e^{-i\omega_{\lambda}t}a_{\lambda}.\end{split}$ (26)
The response of the system to the environment is given by the bath correlation function $\alpha(t-s)=\langle B(t)B^{\dagger}(s)\rangle$. For HEOM we need that the bath correlation function is of the form
$\alpha(t)=\sum_{j=1}^{N_{exp}}G_{j}e^{-W_{j}t},$ (27)
where $G_{j}$ and $W_{j}$ may be complex numbers. For increased readability we perform the next steps for $N_{exp}=1$, but the extension to multiple exponentials is straight forward. The reduced dynamics of the system can be written in the path integral form
$\rho_{t}(q,q^{\prime})$ $=\int\mathrm{d}q_{0}\int\mathrm{d}q_{0}^{\prime}\int\mathcal{D}q\int\mathcal{D}q^{\prime}\,$
×$\times e^{i(S[q]-S[q^{\prime}])}F[q,q^{\prime}]\rho_{0}(q_{0},q_{0}),$ (28)
where $S[q]$ is the action of the closed system and $\rho_{t}(q,q^{\prime})=\bra{q}\rho_{t}\ket{q^{\prime}}$. The influence functional contains all information about the coupling of the open system to the environment. In this case it reads
$F[q,q^{\prime}]=e^{-\int_{0}^{t}\mathrm{d}s\int_{0}^{s}\mathrm{d}s^{\prime}\,(q_{s}-q^{\prime}_{s})\left(\alpha(s-s^{\prime})q_{s^{\prime}}-\alpha^{*}(s-s^{\prime})q^{\prime}_{s^{\prime}}\right)}.$ (29)
We define auxiliary states
$\rho^{(n,m)}(q,q^{\prime},t)=$ $\int\mathrm{d}q_{0}\int\mathrm{d}q_{0}^{\prime}\int\mathcal{D}q\int\mathcal{D}q^{\prime}\,$
×$\times e^{i(S[q]-S[q^{\prime}])}F[q,q^{\prime}]\rho_{0}(q_{0},q_{0})Q_{t}^{n}Q_{t}^{\prime m},$ (30)
where $Q_{t}=\int_{0}^{t}\mathrm{d}s\,\alpha(t-s)q_{s}$ and $Q_{t}^{\prime}=\int_{0}^{t}\mathrm{d}s\,\alpha^{*}(t-s)q^{\prime}_{s}$. Then by directly computing the time derivative of the auxiliary state and by using the exact expression for the open system dynamics together with the special form of the bath correlation function we directly arrive to
$\dot{\rho}^{(n,m)}(t)=$ $-i[H,\rho^{(n,m)}]-(nW+mW^{*})\rho^{(n,m)}$
$+[\rho^{(n+1,m)},q]+[q,\rho^{(n,m+1)}]$
$+nGq\rho_{t}^{(n-1,m)}+mG^{*}\rho^{(n,m-1)}q.$ (31)
The equations presented in the main text are a generalization of the above equations in two respects. First, we allow for multiple exponential terms in the expansion for the bath correlation function, and second we allow for the case $q\neq q^{\dagger}$ [48]. Additionally, it can be shown that
$\rho^{(n,m)}=i^{n}(-i)^{m}\operatorname{Tr}_{env}\left((B^{\dagger}(t))^{m}B(t)^{n}\rho_{tot}\right).$ (32)
Appendix C Derivation of the BBGKY-HEOM equations
In the following we detail the derivation of the BBGKY-HEOM equation. For the sake of readability we again focus on the case of a scalar index, corresponding to a single (global) bath that all particles couple to and a single exponential in the bath correlation function. The situation where each particle in the many-body system couples to its own (local) bath is explained in Appendix D. The extension to a vector index is straight forward. We start from the HEOM equation (II.2) and want to employ the BBGKY hierarchy to reduce the dimension of the system.
Following (11) we define the reduced 2-body auxiliary matrices as
$F_{12}^{(n,m)}=N(N-1)\operatorname{Tr}_{3..N}(\rho^{(n,m)}).$ (33)
Their evolution equation can be derived from the HEOM Eq. (II.2) by tracing over all but $2$ particles and assuming that the density matrix is invariant under particle exchange:
$\begin{split}\dot{F}_{12}^{(n,m)}=&N(N-1)\operatorname{Tr}_{3..N}(\dot{\rho}^{(n,m)}),\\ =&-i[H_{1}+H_{2},F_{12}^{(n,m)}]\\ &-\left[(n-m)i\Delta+(m+n)\kappa\right]F_{12}^{(n,m)}\\ &\quad\!+Gn\left((L_{1}+L_{2})F_{12}^{(n-1,m)}+\operatorname{Tr}_{3}(L_{3}F_{123}^{(n-1,m)})\right)\\ &\quad\!\!+G^{*}m\left(F_{12}^{(n,m-1)}(L_{1}^{\dagger}+L_{2}^{\dagger})+\operatorname{Tr}_{3}(L_{3}^{\dagger}F_{123}^{(n,m-1)})\right)\Bigg{)}\\ &\quad\!+\left[F_{12}^{(n+1,m)},L_{1}^{\dagger}+L_{2}^{\dagger}\right]+\left[L_{1}+L_{2},F_{12}^{(n,m+1)}\right].\end{split}$ (34)
Above we used that operators which are traced over can be cyclically moved under the trace, such that
$\begin{split}\operatorname{Tr}_{3...N}([\sum_{i=1}^{N}H_{i},\rho]=&[H_{1}+H_{2},\rho_{12}]\\ &+\sum_{i=3}^{N}\operatorname{Tr}_{3...N}(H_{i}\rho)-\operatorname{Tr}_{3...N}(\rho H_{i}),\\ =&[H_{1}+H_{2},\rho_{12}].\end{split}$ (35)
In addition, by utilizing the particle exchange symmetry of $\rho^{(n,m)}$, partial traces over operators can be combined with the help of the particle exchange operator $P_{3j}$ for arbitrary $j\in[4,N]$:
$\begin{split}\operatorname{Tr}_{3..N}(L_{j}\rho^{(m-1,p)})=&\operatorname{Tr}_{3..N}(P_{3j}^{2}L_{j}P_{3j}^{2}\rho^{(m-1,p)}),\\ =&\operatorname{Tr}_{3..N}(L_{3}P_{3j}\rho^{(m-1,p)}P_{3j}),\\ =&\operatorname{Tr}_{3..N}(L_{3}\rho^{(m-1,p)}),\\ =&\frac{\operatorname{Tr}_{3}(L_{3}F_{123}^{(m-1,p)})}{N(N-1)(N-2)}.\end{split}$ (36)
In the next step we use a cluster expansion to express the term $F_{123}^{(n,m)}$ in terms of the reduced density one- and two-body matrices. The key approximation that enables this is the neglect of higher order correlations. Combing Eq. (32) and Eq. (11) we can express the 3-body density matrix as
$\begin{split}F_{123}^{(n,m)}=&N(N-1)(N-2)i^{n}(-i)^{m}\\ &\times\operatorname{Tr}_{4..N,env}\left((B^{\dagger}(t))^{m}B(t)^{n}\rho_{tot}\right).\end{split}$ (37)
The cluster expansion amounts to expressing the 3-body density matrix including the environment as:
$\begin{split}\rho_{123E}=&\rho_{1}\otimes\rho_{2}\otimes\rho_{3}\otimes\rho_{E}\\ &+\rho_{1}\otimes\rho_{2}\otimes\Gamma_{3E}+\rho_{1}\otimes\Gamma_{23}\otimes\rho_{E}+\Gamma_{12}\otimes\rho_{3}\otimes\rho_{E}\\ &+\Gamma_{13}\otimes\rho_{2}\otimes\rho_{E}+\Gamma_{1E}\otimes\rho_{2}\otimes\rho_{3}+\rho_{1}\otimes\Gamma_{2E}\otimes\rho_{3}\\ &+\Gamma_{12}\otimes\Gamma_{3E}+\Gamma_{1E}\otimes\Gamma_{23}+\Gamma_{13}\otimes\Gamma_{2E}\\ &+\rho_{1}\otimes\Gamma_{23E}+\Gamma_{13E}\otimes\rho_{2}+\Gamma_{12E}\otimes\rho_{3}+\Gamma_{123}\otimes\rho_{E}\\ &+\Gamma_{1234E}\end{split}$ (38)
We now neglect 3-body correlations within the system $\Gamma_{123}=0$ as well as four body correlations $\Gamma_{123E}=0$ to obtain
$\begin{split}\widetilde{F}^{({n},{m})}_{123}&=4\frac{(N-1)(N-2)}{N^{3}}\operatorname{Tr}(F_{1}^{({n},{m})})F_{1}F_{2}F_{3}\\ &+\frac{N-2}{N}\left(F_{12}F_{3}^{({n},{m})}+F_{2}^{({n},{m})}F_{13}+F_{1}^{({n},{m})}F_{23}\right)\\ &+\frac{N-2}{N}\left(F_{1}F_{23}^{({n},{m})}+F_{13}^{({n},{m})}F_{2}+F_{12}^{({n},{m})}F_{3}\right)\\ &-\frac{N-2}{N^{2}}\operatorname{Tr}(F_{1}^{({n},{m})})\left(F_{12}F_{3}+F_{13}F_{2}+F_{1}F_{23}\right)\\ &-2\frac{(N-2)(N-1)}{N^{2}}\Big{(}F_{1}^{({n},{m})}F_{2}F_{3}\\ &\quad+F_{1}F_{2}^{({n},{m})}F_{3}+F_{1}+F_{2}F_{3}^{({n},{m})}\Big{)}.\end{split}$ (39)
This now leaves us with a reconstruction functional for all auxiliary matrices and enables us to truncate the BBGKY-HEOM equation by neglecting higher order correlations.
Appendix D Local baths
In this section we show how BBGKY-HEOM can be used to treat situations, where each particle couples to its own local bath. For the sake of readability we consider the case where the bath correlation function of each local bath is a single exponential, where, again, the generalization to multiple exponential terms is straight forward. We start out with the HEOM equation for the full N-particle state, where the coupling operators $L_{k}=S_{k}$ only act on particle $k$ and $H_{sys}=\sum_{i=1}^{N}H_{sys}^{i}$
$\begin{split}\dot{\rho}^{(\bm{n},\bm{m})}&=-i[H_{sys},\rho^{(\bm{n},\bm{m})}]-\left(\bm{w}\cdot\bm{n}+\bm{w}^{*}\cdot\bm{m}\right)\rho^{(\bm{n},\bm{m})}\\ &+\sum_{k=1}^{N}\bigg{(}G_{k}\left(n_{k}S_{k}\rho^{(\bm{n}-\bm{e}_{k},\bm{m})}+m_{k}\rho^{(\bm{n},\bm{m}-\bm{e}_{k})}S_{k}^{\dagger}\right)\\ &+\left[\rho^{(\bm{n}+\bm{e}_{k},\bm{m})},S_{k}^{\dagger}\right]+\left[S_{k},\rho^{(\bm{n},\bm{m}+\bm{e}_{k})}\right]\bigg{)},\end{split}$ (40)
Now after tracing out particle $N$, with $\rho_{1..N-1}^{(\bm{n},\bm{m})}=\operatorname{Tr}_{N}(\rho^{(\bm{n},\bm{m})})$ we are left with
$\begin{split}\dot{\rho}_{1..N-1}^{(\bm{n},\bm{m})})&=-i[\sum_{i=1}^{N-1}H_{sys}^{i},\rho_{1..N-1}^{(\bm{n},\bm{m})}]-\left(\bm{w}\cdot\bm{n}+\bm{w}^{*}\cdot\bm{m}\right)\rho_{1..N-1}^{(\bm{n},\bm{m})}\\ &+\sum_{k=1}^{N-1}\bigg{(}G_{k}n_{k}S_{k}\rho^{(\bm{n}-\bm{e}_{k},\bm{m})}_{1..N-1}+G_{k}^{*}m_{k}\rho_{1..N-1}^{(\bm{n},\bm{m}-\bm{e}_{k})}S_{k}^{\dagger}\\ &\quad\quad\quad+\left[\rho_{1..N-1}^{(\bm{n}+\bm{e}_{k},\bm{m})},S_{k}^{\dagger}\right]+\left[S_{k},\rho_{1..N-1}^{(\bm{n},\bm{m}+\bm{e}_{k})}\right]\bigg{)}\\ &+G_{N}n_{N}\operatorname{Tr}_{N}(S_{N}\rho^{(\bm{n}-\bm{e}_{N},\bm{m})})\\ &+G_{N}^{*}m_{N}\operatorname{Tr}_{N}(\rho^{(\bm{n},\bm{m}-\bm{e}_{N})}S_{N}^{\dagger})\Big{)}.\end{split}$ (41)
We see that the last two terms actually still depends on hierarchy states with $n_{N},m_{N}>0$. However, ultimately we are only interested in the physical density matrix $\operatorname{Tr}_{N}(\dot{\rho}^{(\bm{0},\bm{0})})$. We can see that the evolution of this physical density matrix only depends on auxiliary states with $m_{N}=n_{N}=0$, because the sum in line two and three only extends to $N-1$. Furthermore, the evolution of all auxiliary states with $m_{N}=n_{N}=0$ do in turn also not depend on auxiliary states with $m_{N},n_{N}>0$. Thus the physical density matrix is completely independent of this part of the hierarchy and it may as well be omitted. After omitting the last line in Eq. (41) we have thus eliminated the particle $N$ together with its local bath. Continuing to trace out all particles except 1 and 2 we are then left with
$\begin{split}\dot{F}_{12}^{(\bm{n},\bm{m})}&=-i[H_{sys}^{1}+H_{sys}^{2},F_{12}^{(\bm{n},\bm{m})}]-\left(\bm{w}\cdot\bm{n}+\bm{w}^{*}\cdot\bm{m}\right)F_{12}^{(\bm{n},\bm{m})}\\ &+\sum_{k=1}^{2}\bigg{(}G_{k}\left(n_{k}L_{k}F_{12}^{(\bm{n}-\bm{e}_{k},\bm{m})}+m_{k}F_{12}^{(\bm{n},\bm{m}-\bm{e}_{k})}L_{k}^{\dagger}\right)\\ &+\left[F_{12}^{(\bm{n}+\bm{e}_{k},\bm{m})},L_{k}^{\dagger}\right]+\left[L_{k},F_{12}^{(\bm{n},\bm{m}+\bm{e}_{k})}\right]\bigg{)}.\end{split}$ (42)
Appendix E Multipolar gauge
Electric and magnetic fields can be expressed in terms of the vector and scalar potentials
$E$ $=-\nabla\phi-\dot{A},$ (43)
$B$ =×$=\nabla\times A,$ (44)
where $\dot{A}:=\frac{\partial}{\partial t}A$ [98]. The $E$ and $B$ fields are invariant with respect to gauge transformations
$A^{\prime}=A-\nabla\chi,\quad\phi^{\prime}=\phi+\dot{\chi}.$ (45)
We define the Green’s function for the divergence operator as
$\nabla\cdot g^{\parallel}(x,x^{\prime})=-\delta(x-x^{\prime}).$ (46)
We can add to $g^{\parallel}(x,x^{\prime})$ any transversal vector field $g^{\perp}(x,x^{\prime})$ and define
$g_{\mu}(x,x^{\prime})=g^{\parallel}(x,x^{\prime})+\mu g^{\perp}(x,x^{\prime}).$ (47)
The parallel Greens function has the well known representation
$g^{\parallel}(x,x^{\prime})=\nabla\frac{1}{4\pi|x-x^{\prime}|}.$ (48)
A possible choice for the perpendicular Green’s function is
$g^{\perp}(x,x^{\prime})=\int_{0}^{t}\mathrm{d}\lambda x^{\prime}\cdot\delta^{\perp}(\lambda x^{\prime}-x),$ (49)
which corresponds to a straight line from the origin to the point $x^{\prime}$. This particular choice is motivated by the construction of the multipolar gauge [99]. We define a gauge function $\chi_{\mu}$ as
$\chi_{\mu}(x)=\mu\int\mathrm{d}x^{\prime}\,g^{\perp}(x^{\prime},x)A^{\perp}(x^{\prime}).$ (50)
We can thus define new vector potential as
$A_{\mu}(x)=A^{\perp}(x)-\mu\nabla\chi(x).$ (51)
We can define a polarization field also with the help of the Greens function
$P_{\mu}(x)=\int\mathrm{d}x\,g_{\mu}(x,x^{\prime})\rho(x^{\prime}),$ (52)
where $\rho(x)$ is the charge distribution. We see immediately that $\nabla\cdot P_{\mu}(x)=-\rho(x)$. The choice $\mu=0$ corresponds to a Coulomb gauge and $\mu=1$ to the multipolar gauge [99]. We may derive the Hamiltonian in arbitrary $\mu$-gauge from a Lagrangian formalism and we find that [69]
$H=$ $\sum_{i}\frac{1}{2m_{i}}(p_{i}-qA_{\mu}(r_{i}))^{2}+V$
$+\frac{1}{2}\int\mathrm{d}x\,(\Pi_{\mu}(x)+P_{\mu}^{\perp}(x))^{2}+(\nabla\times A^{\perp}(x))^{2}),$ (53)
where $p_{i},x_{i}$ and $A^{\perp},\Pi_{\mu}$ are canonically conjugate and satisfy the commutation relations $[x_{i,a},p_{j,b}]=i\hbar\delta_{ij}\delta_{ab}$ and $[A^{\perp}_{a}(x),\Pi_{\mu,b}(x^{\prime})]=i\hbar\delta_{ab}^{\perp}(x-x^{\prime})$.
We assume that the center of mass of the system is at the origin and do the electric dipole approximation [69]. The Hamiltonian simplifies to
$H$ $=\sum_{i}\frac{(p_{i}-(1-\mu)A_{\mu}(0)^{2})^{2}}{2m}+V$ (54)
$+\sum_{i}\mu d_{i}\cdot\Pi_{\mu}(0)+\frac{1}{2}\mu^{2}\sum_{i}d_{i}\delta^{\perp}(0)d_{i}$ (55)
$+\frac{1}{2}\int\mathrm{d}x\,(\Pi_{\mu}(x)^{2}+(\nabla\times A^{\perp}(x))^{2}).$ (56)
In the Coulomb gauge $(\mu=0)$ we obtain that the coupling to the electromagnetic field is provided by the term $(p_{i}-A_{\mu}(0))^{2}$ leading to the diamagnetic $A^{2}$ term.
On the other hand, in the multipolar gauge ($\mu=1$) the transition dipole of the system $d_{i}$ couples to the field $\Pi_{1}$. Physically this means that the system couples to the field which is the difference of the transverse electric field generated bu the transverse vector potential and the polarization field
$\Pi_{1}=E^{\perp}-P^{\perp}.$ (57)
In addition a self-polarization term arises.
The choice of the gauge leads to different types of interpretation of the physical nature of the electric field the system couples to. We refer further discussions on this subtle points to literature [98, 100, 99, 69]
Appendix F Derivation of the few emitter lasing Hamiltonian
F.1 Huang-Rhys factors from TDDFT
All (TD)DFT calculations have been performed with the ORCA code [101] using the PBE exchange-correlation potential [102] with Grimme D3 dispersion correction [103] and def2-TZVPD basis. We expect the (electronic) structure of methylene-blue (MB) molecules to depend heavily on the embedding in cucurbit[7]uril and in the presence of the nanoparticle-on-mirror set-up. For simplicity, we have assumed here the oxidized form (positively charged MB+) of a single MB molecules, which is typically associated with an intense blue color.
First, the ground-state geometry was converged, and electronic absorption spectra were calculated using the Casida-TDDFT approach (employing the Tamm-Dancoff approximation). Second, the geometry of the S1 and S2 states have been relaxed. We found the S1 state to be virtually dark for absorption and emission (with a Stokes shift of 0.269 eV) and have therefore decided to focus on the S2 transition (Stokes shift 0.119 eV) for subsequent discussion. We note that the selected level of theory is insufficient for direct comparison with experimental spectra as the encapsulation can be expected to play a major role. Using higher-level exchange-correlation potentials (B3LYP+D4) and a continuous polarization model for water flips the excitation order (S1 is then bright) but leads otherwise to minor changes in the excitation frequency (486.8 nm) and oscillator strength at a considerably increased cost. The Tamm-Dancoff approximation has, on the other hand, a notable impact. Full Casida-TDDFT shifts the bright absorption (S1 now) to 550.5 nm and future studies that aim at full theoretical consistency are therefore advised to avoid the Tamm-Dancoff approximation.
In the following we give a brief derivation of Eq. (III.2) and show how the coupling strengths and frequencies of the vibrational modes are obtained from the DFT calculations. In the harmonic approximation the (classical) Hamiltonian governing the dynamics of the nuclear coordinates $\bm{\eta}$, expressed as deviations from their respective equilibrium positions can be given in terms of the mass matrix $\mathbf{M}$ and the Hessian Matrix of the potential energy $\mathbf{K}$ (obtained from the DFT calculations) as
$\begin{split}H=&\frac{1}{2}\dot{\bm{\eta}}^{T}\textbf{M}\dot{\bm{\eta}}+\frac{1}{2}\bm{\eta}^{T}\textbf{K}\bm{\eta},\\ =&\frac{1}{2}\dot{\bm{\mu}}^{2}+\frac{1}{2}\bm{\mu}^{T}\textbf{M}^{-1/2}\textbf{K}\textbf{M}^{-1/2}\bm{\mu},\\ \bm{\mu}=&\textbf{M}^{1/2}\bm{\eta}.\end{split}$ (58)
The mass re-normalized vibrational eigenmodes $u_{k}$ and the corresponding frequencies $\omega_{l}$ are now found by means of diagonalization
$\left[\textbf{M}^{-1/2}\textbf{K}\textbf{M}^{-1/2}\right]\mathbf{u}_{k}=\omega_{k}^{2}\mathbf{u}_{k}.$ (59)
In the electronic groundstate the vibrations are now decoupled in the eigenbasis
$\begin{split}x_{k}=&\bm{\mu}\cdot\mathbf{u}_{k},\\ H=&\frac{1}{2}\sum_{k}\dot{x}_{k}^{2}+\omega_{k}^{2}x_{k}^{2}\end{split}$ (60)
and can be canonically quantized
$\hat{H}=\sum_{k}\frac{\hat{p}_{k}}{2}+\frac{\omega_{k}^{2}}{2}\hat{q}_{k}^{2}.$ (61)
In the double harmonic approximation the dynamics in electronically excited state are governed by the same frequencies but shifted equilibrium positions $\bm{\mu}\to\bm{\mu}-\bm{\Delta}$, where $\bm{\Delta}$ is the mass weighted nuclear displacement between absorption (at 479.5 nm) and emission (at 502.5 nm) $\bm{\Delta}=\textbf{M}^{1/2}(\textbf{R}_{S0}-\textbf{R}_{S2})$. Along the $k$-th eigenmode we then obtain a displacement of $\Delta_{k}=\bm{\Delta}\cdot\textbf{u}_{k}$ in the excited state, leading the the Hamiltonian $H_{mol}$ in Eq. (III.2). The Huang-Rhys factors defined as $S_{k}=\omega_{k}\Delta_{k}^{2}/2$ are thus obtained as
$S_{k}=$ $\frac{\omega_{k}}{2}((\textbf{R}_{S0}-\textbf{R}_{S2})\cdot\sqrt{\textbf{M}}\cdot\textbf{u}_{k})^{2}.$ (62)
F.2 Effective spectral density of intra-molecular vibrations
Eq. (III.2) describes an isolated molecule, where the vibrations form the stick spectrum shown in Fig. 7.
Figure 7: Squared coupling strengths of the vibrational modes (blue) as obtained from DFT and ohmic spectral density of the phonon environment (red).
To account for the environment in which the molecules are embedded we couple each molecular vibration to an (accoustic) phonon bath with an ohmic spectral density. It is important to note that the phonon bath will always relax the vibrational modes to their respective groundstate, which is shifted if the molecule is electronically excited. To ensure this, we model the coupling between the molecular vibrations and the environmental phonon modes with the coupling operator $b_{\lambda}-\Delta_{\lambda}\sigma_{+}\sigma_{-}$ [104]. The total Hamiltonian for a single molecule and the phonon bath with annihilation operators $d_{\xi}$ now reads
$\begin{split}H_{mol,env}=&\frac{\omega_{0}}{2}\sigma^{z}-\frac{1}{\sqrt{2}}\sum_{\lambda=1}^{107}\omega_{\lambda}^{vib}\sqrt{S_{\lambda}}\sigma_{+}\sigma_{-}(b_{\lambda}+b_{\lambda}^{\dagger})\\ &+\sum_{\lambda=1}^{107}\omega_{\lambda}^{vib}b_{\lambda}^{\dagger}b_{\lambda}+\sum_{\lambda,\xi}\omega_{\xi}d_{\lambda,\xi}^{\dagger}d_{\lambda,\xi}\\ &+\sum_{\lambda,\xi}g_{\xi}(b_{\lambda}-\Delta_{\lambda}\sigma_{+}\sigma_{-})d_{\lambda,\xi}^{\dagger}+\mathrm{h.c.},\\ J_{ohm}(\omega)=&\sum_{\xi}\lvert g_{\xi}\rvert^{2}\delta(\omega-\omega_{\xi}),\\ =&A\omega\exp(-\omega/\Lambda)\theta(\omega).\end{split}$ (63)
To achieve a reasonable broadening choose $A=0.025\;\mathrm{cm}^{-2}$ and $\Lambda=3500\;\mathrm{cm}^{-1}$, leading to the spectral density shown in the background of Fig. 7.
Ref. [83] shows how each of the 107 damped vibrational modes can equally be described by an effective spectral density
[…]
We thus broaden each line in Fig. 8 according to Eq. (64), which leads us to the effective spectral density seen by the electronic degree of freedom $J_{eff}(\omega)$ shown in Fig. 3(d) and the effective molecular Hamiltonian
$\begin{split}H_{mol,eff}=&\frac{\omega_{0}}{2}\sigma^{z}-\sum_{\lambda}g_{\lambda}^{eff}\sigma_{+}\sigma_{-}(b_{\lambda}+b_{\lambda}^{\dagger})\\ +&\sum_{\lambda}\omega_{\lambda}^{eff}b_{\lambda}^{\dagger}b_{\lambda},\\ J_{eff}(\omega)=&\sum_{\lambda}\lvert g_{\lambda}^{eff}\rvert^{2}\delta(\omega-\omega_{\lambda}^{eff}).\end{split}$ (65)
The spectral density $J_{eff}(\omega)$ was used to parameterize our BBGKY-HEOM model. In Fig. 8 we compare the resulting absorption spectrum of a single molecule to the path-integral implementation of ORCA [105] and observe qualitative agreement, with some quantitative deviations. A quantitative comparison is complicated by the different treatment of linewidth-broadening in ORCA.
Figure 8: Absorption spectrum of the molecule based on the Hamiltonian (65) (blue) compared to the absorption spectrum predicted by ORCA (dashed black). A quantitative comparison is complicated by the different treatment of linewidth-broadening in ORCA.
F.3 Interaction picture dynamics for few emitter lasing
From the description of a single molecule in Eq. (65) we now move to multiple molecules that couple identically with strength $g^{cav}$ to a common cavity mode that is in resonance with the electronic transition and are subject to a coherent drive with amplitude $2E_{d}$ and frequency $\omega_{0}$
$H_{tot}=$ $\sum_{j}^{N}H_{mol,eff}^{(j)}+\sum_{j}^{N}2E_{d}\cos(\omega_{0}t)\sigma_{x}^{j}$
$+\sum_{j}g^{cav}\sigma_{x}^{j}(a+a^{\dagger})+\omega_{0}a^{\dagger}a,$ (66)
where $H_{mol,eff}^{(j)}$ denotes $H_{mol,eff}$ from Eq. (65) acting on molecule $j$. Cavity loss and spontaneous decay of the individual molecules are accounted for via a master equation description, such that the total system $\rho_{tot}$ consisting of the molecules and the cavity mode evolves according to
$\dot{\rho}_{tot}=-i[H_{tot},\rho_{tot}]+\kappa\mathcal{D}[a](\rho_{tot})+\Gamma_{\downarrow}\sum_{i}\mathcal{D}[\sigma_{-}^{i}](\rho_{tot}).$ (67)
We now go into an interaction picture $\rho_{tot}^{I}=U_{I}\rho U_{I}^{\dagger}$ with
$U_{I}=\exp\left(i\omega_{0}t\left(\sum_{j,\lambda}\sigma_{z}^{j}/2+a_{\lambda}^{\dagger}a_{\lambda}\right)\right).$ (68)
After applying the rotating wave approximation as described in Sec. III.1 we end up with Eq. (III.2),(20).
Appendix G Vibrational Resonances
In the following we provide a heuristic explanation of the maxima observed in Fig. 3. A more detailed analytical investigation based on the full model instead of the toy model ind Eq. (22) is left for future work.
Starting from $H_{simple}$ in Eq. (22) we go into an interaction picture with respect to the drive |$\ket{\widetilde{\psi}}=\exp\left(i\Omega\sigma_{x}+2\Omega b^{\dagger}b\right)\ket{\psi}$ which yields
$\begin{split}\widetilde{H}_{simple}=&(\omega^{vib}-2E_{d})b^{\dagger}b+g^{vib}(\sigma_{+x}b+\sigma_{-x}b^{\dagger})\\ &+\frac{g^{vib}}{2}\left(be^{-2iE_{d}t}+b^{\dagger}e^{2iE_{d}t}\right)\\ &+\frac{g^{vib}}{2}\left(e^{-4iE_{d}t}b^{\dagger}\sigma_{+x}+e^{4iE_{d}t}b\sigma_{-x}\right),\\ \sigma_{+x}=&\frac{\sigma_{z}-i\sigma_{y}}{2},\\ \sigma_{-x}=&\frac{\sigma_{z}+i\sigma_{y}}{2}.\end{split}$ (69)
To obtain an estimate of the phonon population we neglect the time-dependent terms and obtain the dynamics of the expectation value $\langle b\rangle$ in a Heisenberg picture. We find
$\dot{\langle b\rangle}=-i(\omega^{vib}-2E_{d})\langle b\rangle-ig^{vib}\langle\sigma_{-x}\rangle.$ (70)
In the steady state we thus find $\langle b\rangle\propto\frac{1}{\omega^{vib}-2E_{d}}$. As an additional proof of this resonance phenomena is shown in Fig. 9. There we artificially modify the effective spectral density by omitting the small peak at $\omega\approx 1.3E_{max}$. We then find that also the 2nd maxima in the cavity occupation at $E_{d}\approx 0.65E_{max}$ vanishes.
Figure 9: The upper plot shows the obtained steady state cavity occupations for $N=5$ and a modified effective spectral density $J_{eff}(\omega)$, where the small peak at $\omega\approx 1.3E_{max}$ was intentionally omitted as shown in the lower plot. All other parameters are identical to Fig. 3. Compared to the results of the full model the second maxima in the cavity occupation is missing (upper plot).
Appendix H Numerical details
Numerical simulations were performed with the python packages numpy[106], scipy[107] and qutip[108]. The BBGKY-HEOM equation (II.2) with the approximation for the three-body reduced density matrix given in Eq. (II.2) were solved with the scipy implementation of the explicit Runge-Kutta method of order 5(4), with an absolute tolerance of $101e-$ 10 and dynamical timestepping. We checked for convergence with respect to the hierarchy depth, which justifies a hierarchy depth of 5 for the Tavis-Cummings model and 3 for the Cavity-Fermi-Hubbard model. The benchmark simulations for the quantum master equation were performed using qutips mesolve function with default parameters.